# Understanding Mobile App Reviews to Guide Misuse Audits

### v2

@@ I Introduction
Traditional audits of mobile apps conduct a review of their source code (García et al., 2021; Xia et al., 2015). However, interpersonal misuse arising from app users (instead of app developers) goes unnoticed by such processes. We introduce misuse audit, a process for auditing mobile apps through the reported misuse cases.
@@
Research (Chatterjee et al., 2018; Freed et al., 2019; Havron et al., 2019) shows that exploitable apps may cause discomfort, fear, and potential harm to the victim. Possible ways to prevent this risk include highlighting these apps and their exploitable functionalities to users, app distribution platforms, and app developers. However, identifying these apps is nontrivial since they have a legitimate purpose but are misused by abusers. We propose an approach, MissAuditor, for misuse audits that identify exploitable apps and uncover their exploitable functionalities.
@@
Example I shows three reviews (edited for grammar), taken from Apple’s App Store Store (2023a) that are relevant to exploitable behaviors. Although our study is based on Apple App Store’s reviews, MissAuditor can be applied to reviews from other sources, including Google’s Play Store Store (2023b).
@@
(for the AirBeam Video Surveillance app Store (2023c))
@@
In Example I, the first review, for AirBeam Video Store (2023c), addresses the scenario where the app assists a user in accessing a victim’s information without the victim’s knowledge. AirBeam Video is a surveillance app to be installed on the abuser’s device. Hence, the victim may not be an app user but a bystander. The second review, for Life360 Store (2023d), complains about the problem of inappropriate access to the user’s location by the user’s mother. Due to the unequal power dynamics between the victim (reviewer in this case) and the abuser (mother in this case), the victim is forced to install apps that violate their privacy. The third review, from 3Fun Store (2023e), describes a story of improper access to profile pictures. Even though the profile pictures are public, the victim is uncomfortable with the access. It is common for users to upload such information (pictures in this case) on an app with expectations of how other users would access it. As shown in these three cases, information access may lead to discomfort, fear, or potential harm (Freed et al., 2019; Havron et al., 2019). Thus, such cases of information access indicate misuse.
@@ II-A Seed Dataset
Chatterjee et al. Chatterjee et al. (2018) identified 2,707 iOS apps as candidates for Intimate Partner Surveillance (IPS), i.e., apps used by a person to spy on their intimate partner. IPS apps are specific to intimate partners and, hence, a subclass of our concept of exploitable apps. We started our analysis from Chatterjee et al.’s IPS candidate list.
@@ III-A Alarmingness of Reviews
We extracted linguistic features of 1,884 reviews through the Universal Sentence Encoder (USE) Cer et al. (2018), a widely used approach that has proved effective for app reviews (Guo and Singh, 2020; Devine et al., 2021). Using USE features, we trained various multitarget (the targets here being convincingness and severity scores) regression models Borchani et al. (2015) and found that the Support Vector Regressor (SVR) outperforms others. We considered using the reviews’ metadata, such as ratings and titles, but found the metadata to be unhelpful for identifying misuse and left it out. Not only negative ratings and titles but also reviews with positive ratings and titles indicate misuse (partly discussed in Section II-B). For example, an abuser writes a review with the title “Great app” and provides five-star ratings as the app facilitates misuse. As these nondiscriminatory attributes may confuse the model, we decided not to include them while training. We used only the text of reviews for training.
@@ III-B1 Similar Apps
Surveillance apps that can be misused for spying fall under the “Utilities” category, making it an important category to audit. We considered 100 popular utility apps (mentioned on the Apple’s App Store page Store (2023f)) and collected their reviews. Since a popular app can have a lot of reviews (over the years), collecting them can be computationally expensive.
@@ IV RQfunctionality: Uncovering exploitable Functionalities
Some apps monitor a victim’s phone activities, such as browsing history and text messages. Such apps are installed on the victim’s device and activities can be monitored on another synced device. For example, SaferKid Text Monitoring App Store (2023g) allows synced devices to monitor call history, web history, texts, and so on.
@@
Some apps enable audio or video surveillance without the victim’s knowledge. These apps listen, view, or record a victim’s voice or actions and some of them need not be installed on the victim’s phone. For example, Find My Kids: Parental control Store (2023h) is misused to record private conversations between people without them knowing.
@@
Some Global Positioning System (GPS) apps enable tracking a victim’s phone, making them uncomfortable with access. For example, Find My iPhone Store (2023i) is a legitimate app but can be misused to spy on the location of connected devices.
@@
Some apps are misused for stalking user profiles or user-generated content (such as text and images). For example, Kik Messaging and Chat App Store (2023j) can be misused for stalking images and victims’ other information on the app.
Table IV shows these four types of exploitable functionalities and some alarming reviews reporting them. Some of these reviews are old (2014 or 2012), but we confirmed that similar concerns are being raised in the recent reviews of the same apps. For example, the Find My iPhone Store (2023i) app lets its users see the location of the connected devices. Due to unequal power dynamics between the abuser and the victim (say in a family setting) Chatterjee et al. (2018), the victim is sometimes forced to connect to such apps and allow their device to be located. In other words, legitimate apps designed for locating loved ones can be misused for exploitable behavior when the requirements for consent (Singh, 2022) are violated. Hence, they are dual-use Chatterjee et al. (2018). Through a qualitative study, Freed et al. (2018) found that many find-my-phone applications are intended for anti-theft and safety purposes but are heavily misused against privacy. Chatterjee et al. (2018) categorize such apps as spyware, especially in the case of intimate partner surveillance. Just because an app has some legitimate uses doesn’t justify its possible misuse. Hence, we consider such apps to be exploitable. Relying on app reviews highlights both types of exploitable functionalities (always malicious and possibly with legitimate uses) against which future and current app users should be warned. Moreover, app distribution platforms and developers should mitigate privacy risks.
We illustrate the exploitable behavior of the SaferKid Text Monitoring App Store (2023g) by installing it on two devices: a parent’s device (iOS version 14.4.1) and a child’s device (Android version 11.0). Figure 5 shows the exploitable features present in this app. Activities on the child’s device can be monitored on the synced parent’s device. Figure 5(a) shows SaferKid exploitable functionalities such as monitoring text messages, web history, and call history. We verified each of these functionalities. Figure 5(b) shows the screen displaying all chats of the victim. Apps such as SaferKid are advertised as safety apps for children but can be secretly or forcefully installed on another device to monitor the user’s activity. Not only parents but anyone can misuse such apps by installing them on a victim’s phone.
(a) Find My iPhone Store (2023i) app enables the abuser to track multiple devices at a time.
@@
Figure 4: Exploitable functionalities in Find My iPhone Store (2023i). The names and locations of devices are hidden for anonymity.
(a) The SaferKid Store (2023g) app provides multiple ways to monitor victim’s phone activities.
(b) Using the SaferKid Store (2023g) app, victim’s chats can be seen on the abuser’s phone.
Figure 5: Exploitable functionalities in SaferKid Text Monitoring Store (2023g). The actual chats and device’s name are hidden for anonymity.
@@
Monitoring phone activities SaferKid Text Monitoring App Store (2023g) …Tracking things like social media, texts, and search history is just a complete disregard of privacy. You have to have trust in your kids …Apps like these shouldn’t be allowed. IF YOU TRUST YOUR KID, DONT DOWNLOAD. (Date: 2019-12-07)
Audio or video surveillance Find My Kids: Parental control Store (2023h) This app proves to have a invasion of privacy. Due to the fact if your kids was at a friends house and talking to his friends parents, this app records what is going on and is a invasion of privacy. If your child left their phone downstairs or anywhere and they are playing it can record private conversation between adult and is a unsafe … (Date: 2019-01-16)
Tracking location Find My iPhone Store (2023i) …It’s supposed to be used to recover a lost phone, not to religiously stalk your children.…The fact that a mom actually installed this app onto her son’s phone without his knowledge is flat out wrong. …If you’re constantly monitoring your child 24/7, just imagine what your child will do when they go off to college. … (Date: 2014-02-13)
Profile stalking Kik Messaging and Chat App Store (2023j) MAKE THIS SAFE PEOPLE WANT TO USE IT BUT DON’T WANT TO BE STALKED OR ABUSED IN THE APP PROTECT THE APP OR PEOPLE WON’T USE IT MY FRIEND HAD A MAN SEND HER A BAD PICTURE IF YOU KNOW WHAT I MEAN. OVERALL THIS APP IS GREAT!!!!!” (Date: 2015-06-24)
@@ V-A Auditing and App Reviews
Most research on software audits analyzes the flow and dependencies in source code (García et al., 2021; Xia et al., 2015; Perl et al., 2015; Pashchenko et al., 2020). Only a few audit studies consider data such as system traces Wang et al. (2015). However, all these studies fail to identify exploitable apps because they focus on the technical aspects of building or running them. Thus, their analysis doesn’t address cases of forced consent and inappropriate access to public information. Our work leverages app reviews, unexplored by prior studies, for misuse audits.
@@ V-B Spying through Mobile Apps
Chatterjee et al. Chatterjee et al. (2018) apply search queries and manual verification based on information such as app descriptions and permissions to identify Intimate Partner Surveillance apps. However, in general, the actual usage deviates from the legitimate purpose shown in app descriptions. To identify such misuse, we focus on the evidence provided in app reviews.
Roundy et al. Roundy et al. (2020) identify apps used for phone number spoofing and message bombing, which lie outside our scope of exploitable apps. Conversely, exploitable apps include those that enable stalking public information, which lie outside their scope. Roundy et al. use installation data, to uncover spying apps that are installed on infected devices. However, we focus on evidence of exploitable behavior present in app reviews to uncover exploitable apps. Roundy et al. rely upon Norton’s security app Store (2023k) to determine which devices are infected. Thus, their approach would miss apps that a general user can leverage to spy.
All these studies, along with a few others (Tseng et al., 2020; Zou et al., 2021; Tseng et al., 2021), investigate how technology is abused for spying. However, they do not support the broader set of exploitable apps. In particular, they do not consider cases when the victim is uncomfortable with the access (e.g., of public information) even if aware of it, along with cases involving family, not only strangers, as abusers.
Some studies address user privacy in the context of user information shared with software developers Reardon et al. (2019); Block et al. (2017); Marforio et al. (2012) . They do not consider cases where an app user can exploit the app to access the information of a victim (another app user or bystander).
@@ VI-B2 Threats Remaining
Now, we describe the threats that remain in our work. First, we investigated only a few thousand apps, which may not represent all apps on the Apple App Store Store (2023a). The performance of MissAuditor may be reduced while testing it on all apps of the Apple App Store. Second, we targeted apps and their reviews only on Apple’s App Store. Upon deployment on other app stores, the performance of our approach can differ. Third, suppose an app distribution platform does not offer recommendations for similar apps. In that case, finding good candidates for exploitable apps will be a challenge. An alternative is to prioritize applying MissAuditor for the apps flagged by app users (victims). Fourth, some reviews in our datasets are old. Due to frequent app updates, the types of misuse may change over time. Although we confirm that our findings are still relevant in Section III-B3, there is a possibility that a few apps may not be exploitable anymore. Fifth, some findings are based on a random review sample because manually reading all reviews is not feasible. The results may vary in the latter case.
@@ VI-C Limitations and Future Directions
Fifth, app reviews can be ambiguous in differentiating between inherently malicious and benign but exploitable functionalities. For example, the Safer Kid Text Monitoring App Store (2023g) is designed to keep loved ones safe (benign app) but is misused according to reviews. Future research can distinguish between these benign apps and other malicious ones.
@@ VII Reproducibility
We have released our entire training data along with the computational model on Zenodo (Garg et al., 2024). Our appendix provides details of (1) dataset collection and annotation, (2) training and evaluation, and (3) linguistic analysis.

### v3

@@ I Introduction
Traditional audits of mobile apps conduct a review of their source code [14, 52]. However, interpersonal misuse arising from app users (instead of app developers) goes unnoticed by such processes. We introduce misuse audit, a process for auditing mobile apps through the reported misuse cases.
@@
Research [6, 13, 19] shows that exploitable apps may cause discomfort, fear, and potential harm to the victim. Possible ways to prevent this risk include highlighting these apps and their exploitable functionalities to users, app distribution platforms, and app developers. However, identifying these apps is nontrivial since they have a legitimate purpose but are misused by abusers. We propose an approach, MissAuditor, for misuse audits that identify exploitable apps and uncover their exploitable functionalities.
@@
Example I shows three reviews (edited for grammar), taken from Apple’s App Store [37] that are relevant to exploitable behaviors. Although our study is based on Apple App Store’s reviews, MissAuditor can be applied to reviews from other sources, including Google’s Play Store [47].
@@
(for the AirBeam Video Surveillance app [36])
@@
In Example I, the first review, for AirBeam Video [36], addresses the scenario where the app assists a user in accessing a victim’s information without the victim’s knowledge. AirBeam Video is a surveillance app to be installed on the abuser’s device. Hence, the victim may not be an app user but a bystander. The second review, for Life360 [41], complains about the problem of inappropriate access to the user’s location by the user’s mother. Due to the unequal power dynamics between the victim (reviewer in this case) and the abuser (mother in this case), the victim is forced to install apps that violate their privacy. The third review, from 3Fun [45], describes a story of improper access to profile pictures. Even though the profile pictures are public, the victim is uncomfortable with the access. It is common for users to upload such information (pictures in this case) on an app with expectations of how other users would access it. As shown in these three cases, information access may lead to discomfort, fear, or potential harm [13, 19]. Thus, such cases of information access indicate misuse.
@@ II-A Seed Dataset
Chatterjee et al. [6] identified 2,707 iOS apps as candidates for Intimate Partner Surveillance (IPS), i.e., apps used by a person to spy on their intimate partner. IPS apps are specific to intimate partners and, hence, a subclass of our concept of exploitable apps. We started our analysis from Chatterjee et al.’s IPS candidate list.
@@ III-A Alarmingness of Reviews
We extracted linguistic features of 1,884 reviews through the Universal Sentence Encoder (USE) [5], a widely used approach that has proved effective for app reviews [16, 9]. Using USE features, we trained various multitarget (the targets here being convincingness and severity scores) regression models [4] and found that the Support Vector Regressor (SVR) outperforms others. We considered using the reviews’ metadata, such as ratings and titles, but found the metadata to be unhelpful for identifying misuse and left it out. Not only negative ratings and titles but also reviews with positive ratings and titles indicate misuse (partly discussed in Section II-B). For example, an abuser writes a review with the title “Great app” and provides five-star ratings as the app facilitates misuse. As these nondiscriminatory attributes may confuse the model, we decided not to include them while training. We used only the text of reviews for training.
@@ III-B1 Similar Apps
Surveillance apps that can be misused for spying fall under the “Utilities” category, making it an important category to audit. We considered 100 popular utility apps (mentioned on the Apple’s App Store page [46]) and collected their reviews. Since a popular app can have a lot of reviews (over the years), collecting them can be computationally expensive.
@@ IV RQfunctionality: Uncovering exploitable Functionalities
Some apps monitor a victim’s phone activities, such as browsing history and text messages. Such apps are installed on the victim’s device and activities can be monitored on another synced device. For example, SaferKid Text Monitoring App [44] allows synced devices to monitor call history, web history, texts, and so on.
@@
Some apps enable audio or video surveillance without the victim’s knowledge. These apps listen, view, or record a victim’s voice or actions and some of them need not be installed on the victim’s phone. For example, Find My Kids: Parental control [38] is misused to record private conversations between people without them knowing.
@@
Some Global Positioning System (GPS) apps enable tracking a victim’s phone, making them uncomfortable with access. For example, Find My iPhone [39] is a legitimate app but can be misused to spy on the location of connected devices.
@@
Some apps are misused for stalking user profiles or user-generated content (such as text and images). For example, Kik Messaging and Chat App [40] can be misused for stalking images and victims’ other information on the app.
Table IV shows these four types of exploitable functionalities and some alarming reviews reporting them. Some of these reviews are old (2014 or 2012), but we confirmed that similar concerns are being raised in the recent reviews of the same apps. For example, the Find My iPhone [39] app lets its users see the location of the connected devices. Due to unequal power dynamics between the abuser and the victim (say in a family setting) [6], the victim is sometimes forced to connect to such apps and allow their device to be located. In other words, legitimate apps designed for locating loved ones can be misused for exploitable behavior when the requirements for consent [34] are violated. Hence, they are dual-use [6]. Through a qualitative study, Freed et al. [12] found that many find-my-phone applications are intended for anti-theft and safety purposes but are heavily misused against privacy. Chatterjee et al. [6] categorize such apps as spyware, especially in the case of intimate partner surveillance. Just because an app has some legitimate uses doesn’t justify its possible misuse. Hence, we consider such apps to be exploitable. Relying on app reviews highlights both types of exploitable functionalities (always malicious and possibly with legitimate uses) against which future and current app users should be warned. Moreover, app distribution platforms and developers should mitigate privacy risks.
We illustrate the exploitable behavior of the SaferKid Text Monitoring App [44] by installing it on two devices: a parent’s device (iOS version 14.4.1) and a child’s device (Android version 11.0). Figure 5 shows the exploitable features present in this app. Activities on the child’s device can be monitored on the synced parent’s device. Figure 5(a) shows SaferKid exploitable functionalities such as monitoring text messages, web history, and call history. We verified each of these functionalities. Figure 5(b) shows the screen displaying all chats of the victim. Apps such as SaferKid are advertised as safety apps for children but can be secretly or forcefully installed on another device to monitor the user’s activity. Not only parents but anyone can misuse such apps by installing them on a victim’s phone.
(a) Find My iPhone [39] app enables the abuser to track multiple devices at a time.
@@
Figure 4: Exploitable functionalities in Find My iPhone [39]. The names and locations of devices are hidden for anonymity.
(a) The SaferKid [44] app provides multiple ways to monitor victim’s phone activities.
(b) Using the SaferKid [44] app, victim’s chats can be seen on the abuser’s phone.
Figure 5: Exploitable functionalities in SaferKid Text Monitoring [44]. The actual chats and device’s name are hidden for anonymity.
@@
Monitoring phone activities SaferKid Text Monitoring App [44] …Tracking things like social media, texts, and search history is just a complete disregard of privacy. You have to have trust in your kids …Apps like these shouldn’t be allowed. IF YOU TRUST YOUR KID, DONT DOWNLOAD. (Date: 2019-12-07)
Audio or video surveillance Find My Kids: Parental control [38] This app proves to have a invasion of privacy. Due to the fact if your kids was at a friends house and talking to his friends parents, this app records what is going on and is a invasion of privacy. If your child left their phone downstairs or anywhere and they are playing it can record private conversation between adult and is a unsafe … (Date: 2019-01-16)
Tracking location Find My iPhone [39] …It’s supposed to be used to recover a lost phone, not to religiously stalk your children.…The fact that a mom actually installed this app onto her son’s phone without his knowledge is flat out wrong. …If you’re constantly monitoring your child 24/7, just imagine what your child will do when they go off to college. … (Date: 2014-02-13)
Profile stalking Kik Messaging and Chat App [40] MAKE THIS SAFE PEOPLE WANT TO USE IT BUT DON’T WANT TO BE STALKED OR ABUSED IN THE APP PROTECT THE APP OR PEOPLE WON’T USE IT MY FRIEND HAD A MAN SEND HER A BAD PICTURE IF YOU KNOW WHAT I MEAN. OVERALL THIS APP IS GREAT!!!!!” (Date: 2015-06-24)
@@ V-A Auditing and App Reviews
Most research on software audits analyzes the flow and dependencies in source code [14, 52, 29, 27]. Only a few audit studies consider data such as system traces [51]. However, all these studies fail to identify exploitable apps because they focus on the technical aspects of building or running them. Thus, their analysis doesn’t address cases of forced consent and inappropriate access to public information. Our work leverages app reviews, unexplored by prior studies, for misuse audits.
@@ V-B Spying through Mobile Apps
Chatterjee et al. [6] apply search queries and manual verification based on information such as app descriptions and permissions to identify Intimate Partner Surveillance apps. However, in general, the actual usage deviates from the legitimate purpose shown in app descriptions. To identify such misuse, we focus on the evidence provided in app reviews.
Roundy et al. [33] identify apps used for phone number spoofing and message bombing, which lie outside our scope of exploitable apps. Conversely, exploitable apps include those that enable stalking public information, which lie outside their scope. Roundy et al. use installation data, to uncover spying apps that are installed on infected devices. However, we focus on evidence of exploitable behavior present in app reviews to uncover exploitable apps. Roundy et al. rely upon Norton’s security app [42] to determine which devices are infected. Thus, their approach would miss apps that a general user can leverage to spy.
All these studies, along with a few others [48, 54, 49], investigate how technology is abused for spying. However, they do not support the broader set of exploitable apps. In particular, they do not consider cases when the victim is uncomfortable with the access (e.g., of public information) even if aware of it, along with cases involving family, not only strangers, as abusers.
Some studies address user privacy in the context of user information shared with software developers [32, 3, 22] . They do not consider cases where an app user can exploit the app to access the information of a victim (another app user or bystander).
@@ VI-B2 Threats Remaining
Now, we describe the threats that remain in our work. First, we investigated only a few thousand apps, which may not represent all apps on the Apple App Store [37]. The performance of MissAuditor may be reduced while testing it on all apps of the Apple App Store. Second, we targeted apps and their reviews only on Apple’s App Store. Upon deployment on other app stores, the performance of our approach can differ. Third, suppose an app distribution platform does not offer recommendations for similar apps. In that case, finding good candidates for exploitable apps will be a challenge. An alternative is to prioritize applying MissAuditor for the apps flagged by app users (victims). Fourth, some reviews in our datasets are old. Due to frequent app updates, the types of misuse may change over time. Although we confirm that our findings are still relevant in Section III-B3, there is a possibility that a few apps may not be exploitable anymore. Fifth, some findings are based on a random review sample because manually reading all reviews is not feasible. The results may vary in the latter case.
@@ VI-C Limitations and Future Directions
Fifth, app reviews can be ambiguous in differentiating between inherently malicious and benign but exploitable functionalities. For example, the Safer Kid Text Monitoring App [44] is designed to keep loved ones safe (benign app) but is misused according to reviews. Future research can distinguish between these benign apps and other malicious ones.
@@ VII Reproducibility
We have released our entire training data along with the computational model on Zenodo [15]. Our appendix provides details of (1) dataset collection and annotation, (2) training and evaluation, and (3) linguistic analysis.